After engineers have continued to monitor the affected switch, there have been no further issues and all services are showing restored with normal traffic volumes.
We are now comfortable this issue has been resolved.
Switch connectivity has been restored and we are monitoring to ensure service stability.
We are currently experiencing a fault with a switch at our 220Q PoP which will be affecting some web hosting, virtual server, dedicated server and colocation customers.
An engineer is en-route to the facility to diagnose.